OpenStudy (anonymous):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned? 44 40 76 64

OpenStudy (jdoe0001):

we know that there are 100 men and women, we could say that m + w = 100 "twice the number of men" => 2m " less " => - "three times the number of women" => 3w "is equal to 20" => = 20 -------------------------------------------- m + w = 100 2m - 3w = 20 ------------------------------- use the substitution method solve the 1st one for either and then substitute it in the 2nd one